# OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis

### v2

@@ 3.2 Novel View Synthesis
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure  3. Example camera interpolated images can be seen in Figure  3, and example images using QANV can be seen in Figure  3
@@ 3.3 OmniAD
Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level. Although Reverse Distillation (RD) [14] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
Backbone Exploration. We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints. Many alternatives struggled to converge or offered minimal gains, but ResNeXt [38] consistently yielded better multi-scale representations and improved performance compared to simpler encoders. In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder. However, given the ever-growing variety of such networks, we restricted our exploration to widely adopted backbones and found that ResNeXt provided the most consistent gains for our multi-view setting.
Receptive Field Constraints. Further analysis indicated that RD’s student decoder had a relatively small effective receptive field (ERF) [15], causing it to overlook anomalies. This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
Rationale of ERF Expansion for Anomaly Sensitivity. Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image. Suppose a teacher network $T$ extracts multi-scale features
$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$
where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map. A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$. In Reverse Distillation [14], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous (normal) data alone.
@@
$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$
@@
Impact of ERF Expansion. Let $A\subset\mathcal{I}$ denote the anomaly region, i.e., pixels in $x$ that deviate from normal. By adding student attention modules (Fig. 6), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$:
$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$
@@
Reverse Distillation Anomaly Score. As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring discrepancy between $\hat{x}_{ij}$ and $x_{ij}$. When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity. With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
Attention Modules for the Student Decoder. To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block. These modules expand the ERF by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies. Figure 6 shows the basic layout of each attention module, incorporating interleaved 3x3 convolutions and self-attention blocks for both local and global feature integration.
Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder. We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
Figure 6: Student attention module.
@@ 4 ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure  8. Each setup also contains different anomalies, as shown in Figure  8. Additionally, each anomaly image has at most one anomaly but each setup has varying anomalies that were added and removed. The dataset includes the manually annotated ground truth for anomaly segmentation. Table 1 provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table 2. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
Figure 8: ToyCity consists of four configurations as shown here. We present an example baseline image for each as well as an example of one of several anomalies used. A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
@@
Name Non-Anomalous Anomalous QANV IBV
ToyCity 1 316 319 319 3781
ToyCity 2 341 308 308 3773
ToyCity 3 307 317 317 3673
ToyCity 4 314 314 314 3757
@@ 5 Experiments
We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
@@ 5.1 Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$ (1)
The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous. This approach would yield numerous true negatives and few false positives. A similar effect can be observed with the AUROC, which may show a more optimistic estimate of model performance. However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives. Thus we provide both the $F_{1}$ score and AUROC.
5.2 Implementation Details
To allow for fair comparison between all anomaly detection methods we use the anomalib [39] library implementation of each anomaly detection method. A fixed random seed of 42 was used for all methods. All methods were allowed to train for a maximum of 100 epochs and the max achieved $F_{1}$ score was selected as the best model. All methods converge well before 100 epochs were completed ensuring fair training. Additionally methods were trained with the same batch size. Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
5.3 Ablation Studies
We conducted ablation studies to validate both the OmniAD architecture and the proposed NVS augmentation strategy.
5.3.1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables 3 and 4 and MAD Tables 7 we can see a consistent performance increase when $A^{i}$ is used.
5.3.2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table 2.
@@ 5.4 Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables 5,6,8. Additional experiments on different scenes is provided in the supplemental material.
The RD with ResNext model and the canonical RD methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model. The shift from v1 to v2, where anomaly aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies. Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ Scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations. This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and RD.
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table 5,6 and MAD Table 8. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
Table 2: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
Variation Non-Anomalous Anomalous QANV INV
v1 - -
v2 -
v3 -
v4
@@ 5.5 Qualitative Evaluation
Figure  9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
Figure 9: Segmentation performance comparison between OmniAD and RD with ResNeXt backbone across four ToyCity anomalies and 2 objects from MAD.We illustrating three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4m and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
Figure 10: Additional segmentations and anomaly scorces from OmniAD.
@@
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
@@
OmniAD-R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
@@
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
@@
OmniAD-R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
@@
OmniAD-R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
OmniAD-$A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
RD 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
CFLOW 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
FastFlow 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
CFA 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028

### v3

@@ 3.2 Novel View Synthesis
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure 4. Example camera interpolated images can be seen in Figure 4, and example images using QANV can be seen in Figure 4
@@ 3.3 OmniAD
Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level. Although Reverse Distillation[25] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios. As such, OmniAD is proposed to refine Reverse Distillation to improve its generalizability in Scene AD, an overview of OmniAD is shown in 6.
Figure 6: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder. We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$.
Backbone Exploration. We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints. Many alternatives struggled to converge or offered minimal gains, but ResNeXt[50] consistently yielded better multiscale representations and improved performance compared to simpler encoders. In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder. However, given the ever-growing variety of such networks, we restricted our exploration to widely adopted backbones and found that ResNeXt provided the most consistent gains for our multi-view setting.
Receptive Field Constraints. Further analysis indicated that Reverse Distillation’s student decoder had a relatively small effective receptive field (ERF) [28], causing it to overlook anomalies. This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
Rationale of ERF Expansion for Anomaly Sensitivity. Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image. Suppose a teacher network $T$ extracts multiscale features,
$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$ (1)
where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map. A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$. In RD [25], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous data alone $S$ is unable to reconstruct anomalous features from a teacher.
@@
$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$ (2)
@@
Impact of ERF Expansion. Let $A\subset\mathcal{I}$ denote the anomaly region, i.e., pixels in $x$ that deviate from normal. By adding student attention modules (Fig. 7), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$,
$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$ (3)
@@
Effect of ERF on Anomaly Score. As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring the discrepancy between $\hat{x}_{ij}$ and $x_{ij}$. When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity. With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
Expansion of ERF with Student Attention Modules.
To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block. These modules expand the ERF, as shown in Figure 8, by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies. Figure 7 shows the basic layout of each attention module and relevant parameters are shown in Table 1.
Table 1: Student attention module parameters and layers: $A^{3}$: $C_{in}=256,H_{in}=64,C_{out}=256,H_{out}=64$. $A^{2}$: $C_{in}=512,H_{in}=32,C_{out}=512,H_{out}=32$. $A^{1}$: $C_{in}=1024,H_{in}=16,C_{out}=1024,H_{out}=16$. Batch Normalization (BN), 2D Max Pooling (MaxPool2D), 2D Convolution (Conv2D)
Stage Layer Type Details Output Shape
Input - Input Channels: $C_{in}$, Spatial Size: $H_{in}\times H_{in}$ $B\times C_{in}\times H_{in}\times H_{in}$
Encoder Block 1 Conv2D kernel=3, stride=1, padding=1, $C_{in}\to 64$ $B\times 64\times H_{in}\times H_{in}$
Self Attention channels=64 $B\times 64\times H_{in}\times H_{in}$
MaxPool2D kernel=2, stride=2 $B\times 64\times H_{in}/2\times H_{in}/2$
Encoder Block 2 Conv2D kernel=3, stride=1, padding=1, $64\to 128$ $B\times 128\times H_{in}/2\times H_{in}/2$
Self Attention channels=128 $B\times 128\times H_{in}/2\times H_{in}/2$
MaxPool2D kernel=2, stride=2 $B\times 128\times H_{in}/4\times H_{in}/4$
Bottleneck Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $128\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $256\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $256\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
Conv2D, BN kernel=3, stride=1, padding=1, $256\to 128$ $B\times 128\times H_{in}/4\times H_{in}/4$
Decoder Block 1 Upsample Target: $H_{in}/2\times H_{in}/2$ $B\times 128\times H_{in}/2\times H_{in}/2$
Conv2D kernel=3, stride=1, padding=1, $256\to 64$ $B\times 64\times H_{in}/2\times H_{in}/2$
Decoder Block 2 Upsample Target: $H_{in}\times H_{in}$ $B\times 64\times H_{in}\times H_{in}$
Conv2D kernel=3, stride=1, padding=1, $128\to C_{out}$ $B\times C_{out}\times H_{in}\times H_{in}$
Figure 7: Student attention module layout.
@@ 4 ToyCity Dataset
The ToyCity dataset addresses the need for a real image benchmark dataset consisting of multiple objects and multiple views. As shown in Table 2, ToyCity is comparable in size to existing prominent AD benchmarks that consist of real images of anomalies. Additionally, ToyCity consists of substantially more real images, consists of unconstrained random views, and is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure 9. Each setup also contains different anomalies, as shown in Figure 10. Additionally, each anomaly image has at most one anomaly, but each setup has varying anomalies that were added and removed as shown in Figure 10. Anomalous images contain on average $0.22\%$ anomalous pixels (standard deviation $0.13\%$). The percentage of anomalous pixels per image ranges from a minimum of $0.03\%$ to a maximum of $0.78\%$. Table 3 provides an overview of the number of images released for ToyCity. Each ToyCity has four augmented variations, without any augmentations, with Query-Aligned Novel Views (QANV), with Interpolated Non-Anomalous Views (INV), and combining both QANV and INV; these variations are presented in Table 4. ToyCity thus offers 16 total variations that will be evaluated.
Table 2: Comparison of datasets containing real images of anomalies with ground truth labels for quantitative evaluation.
Dataset Data Availability Labeled Anomalous Images Multi-View Multi-Object
AITEX[14] Yes 105 No No
MPDD[20] Yes 282 No Yes
MVTec-AD[13] Yes 1258 No No
MVTec 3D-AD[27] Yes 948 No No
MVTec LOCO-AD[16] Yes 993 No Yes
NanoTwice[26] Yes 40 No No
VisA[15] Yes 1200 No Yes
MAD-Real[12] Partial 221 Yes No
ToyCity Yes 1258 Yes Yes
Figure 9: ToyCity consists of four configurations as shown here. We present an example baseline image for each, as well as an example of one of several anomalies used. A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
Figure 10: Each ToyCity configuration has several anomalies that have been added and removed from the scene. Examples of these anomalies for each ToyCity are shown here by a red bounding box.
@@
Name Non-Anomalous Anomalous QANV IBV Images
ToyCity 1 316 319 319 3781 4735
ToyCity 2 341 308 308 3773 4730
ToyCity 3 307 317 317 3673 4614
ToyCity 4 314 314 314 3757 4699
@@ 5 Experiments
We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method. As the camera poses used in MAD-Real [12] for testing were not released publicly, we could only utilize a subset of the MAD dataset that COLMAP could generate camera poses for. Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. Additionally, we evaluate the performance of OmniAD on the MVTec-AD dataset [13]. An ablation study follows to show the significance of the student attention modules on Reverse Distillation. Finally, a qualitative comparison between OmniAD and Reverse Distillation for anomaly segmentation is presented.
@@ 5.1 Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$,
$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$ (4)
The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  9, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous. This approach would yield numerous true negatives and few false positives. A similar effect can be observed with the AUROC, which may show a more optimistic estimate of model performance. However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives. Thus, we provide both the $F_{1}$ score and AUROC.
5.2 Implementation and Training Details
To ensure a fair comparison and promote reproducibility, all anomaly detection methods evaluated in this study (including CFLOW-AD, FastFlow, CFA, DRÆM, RD, and our proposed OmniAD variants) were implemented using the publicly available anomalib library [37], which has faithfully integrated each method based on original implementations and allows for reproducibility.
Parameter Settings: For all baseline methods, we utilized the default hyperparameter configurations provided within the anomalib library to match original implementations. OmniAD and its ablated versions also inherit these same default hyperparameters from the anomalib implementation of the underlying Reverse Distillation method. This approach ensures that performance differences observed are primarily attributable to the architectural modifications (i.e., the ResNeXt backbone and the student attention modules in OmniAD) and the novel view synthesis data augmentation strategies, rather than method-specific hyperparameter tuning.
Training Procedure: All models were trained for a maximum of 100 epochs, using a batch size of 32 for fair comparison. Early stopping was implicitly handled by selecting the model checkpoint that yielded the maximum pixel-level $F_{1}$ score on a validation set, consistent with standard practices. We observed that all methods converged well before the 100-epoch limit, ensuring adequate training time. Learning rates and model specific training hyperparameters follow the anomalib defaults for each model.
@@ 5.3 Quantitative Evaluation
In our assessment of diverse AD methods for Scene AD using ToyCity and MAD-Real[12], Table 5,6,9, we observe a consistent performance increase for OmniAD compared to Reverse Distillation. A similar trend can also be seen for AUROC Table 7,8,10.
The Reverse Distillation with ResNext model and the canonical Reverse Distillation methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model. The shift from no augmentation to QANV, where query aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies due to augmentation. Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations. This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and Reverse Distillation.
As mentioned, metrics such as pixel-level AUROC can often show very high numbers for datasets, such as ToyCity and MAD, due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table 7,8 and MAD Table 10. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
To evaluate the generalizability of OmniAD beyond Scene AD, we also test its performance in a standard fixed-view AD setting using the MVTec-AD dataset[13]. As indicated in Table 11, OmniAD remains effective in this fixed-view context. Although the high AUROC (0.975) achieved by Reverse Distillation for fixed-view AD limits newer methods like OmniAD to incremental improvements until a perfect AUROC is attained. OmniAD does demonstrate a more substantial performance gain in terms of the $F_{1}$ score compared to Reverse Distillation and other recent methods such as SimpleNet[51], MambaAD[34], and AnomalyDINO[52].
Table 4: Four augmentation variations are tested for each of the scenes present in ToyCity.
Variation Non-Anomalous QANV INV Images
No Augmentation - - 300+
QANV Augmentation - 600+
INV Augmentation - 3600+
Both QANV and INV 3900+
@@ 5.4 Qualitative Evaluation
Figure 11 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, Reverse Distillation, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, Reverse Distillation may erroneously identify arbitrary areas as anomalous. Secondly, Reverse Distillation has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, Reverse Distillation often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in Reverse Distillation can also be seen from the MAD dataset, in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD and Reverse Distillation with ResNeXt for each dataset are presented in Figure 12, showing a heat map of the anomaly score as well as the segmentation. Additional comparisons between evaluated methods across different anomalies are shown in Figure 13.
Figure 11: Segmentation performance comparison between the top two evaluated AD methods OmniAD and Reverse Distillation with a ResNeXt backbone across four ToyCity anomalies and objects from MAD. We illustrate three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4, and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig). Other evaluated methods also fail in the same manner as Reverse Distillation with a ResNeXt, but at a higher frequency.
Figure 12: Additional segmentation and anomaly scores from OmniAD.
Figure 13: Additional segmentations across all evaluated AD methods.
@@
OmniAD w/o R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
OmniAD w/o $A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
RD [25] 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
CFLOW-AD [21] 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
FastFlow [22] 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
CFA [23] 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
DRÆM [24] 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
@@
OmniAD w/o R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD w/o $A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD [25] 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW-AD [21] 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow [22] 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA [23] 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM [24] 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
@@
OmniAD w/o R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
OmniAD w/o $A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
RD [25] 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
CFLOW-AD [21] 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
FastFlow [22] 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
CFA [23] 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
DRÆM [24] 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
@@
OmniAD w/o R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD w/o $A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD [25] 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW-AD [21] 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow [22] 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA [23] 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM [24] 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
@@
OmniAD w/o R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
OmniAD w/o $A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
RD [25] 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
CFLOW-AD [21] 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
FastFlow [22] 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
CFA [23] 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
DRÆM [24] 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028